lexxii:

find the geometric mean of 4 and 12

Shadow:

\[GM = \sqrt ( 4 \times 12)\]

lexxii:

I got \[4\sqrt{3}\] is that right? I wasn't sure on this one....

Shadow:

Well, the geometric mean of two numbers means that you multiply those two numbers together, then take the square root. (For three numbers you would take the 3rd root). But you did it correctly.

Shadow:

Just be sure to check what form they want it in.

lexxii:

okay :) thank you!